OpenStudy (anonymous):

A peregrine falcon flew 16 and 2/3 miles in 5 minutes. What is the speed in feet per second? Round your answer to the nearest tenth. Please show your work

OpenStudy (jdoe0001):

keeping in mind that \(\bf 16\cfrac{2}{3}\implies \cfrac{16\cdot 3+2}{3}\implies \cfrac{50}{3}\ and that there are 5280 feet in 1 mile, and 60 secs in 1 minute \(\bf \cfrac{\frac{50}{3}mile}{5min}\implies \cfrac{50}{3}mile\cdot \cfrac{1}{5min}\implies \cfrac{50mile}{15min}\qquad thus\\ \quad \\ \cfrac{50\cancel{mile}}{15\cancel{min}}\cdot \cfrac{5280ft}{\cancel{mile}}\cdot \cfrac{\cancel{min}}{60secs}\)

OpenStudy (jdoe0001):

ohh... shoot... one sec

OpenStudy (jdoe0001):

lemme redo that

OpenStudy (jdoe0001):

keeping in mind that \(\bf 16\cfrac{2}{3}\implies \cfrac{16\cdot 3+2}{3}\implies \cfrac{50}{3}\) and that there are 5280 feet in 1 mile, and 60 secs in 1 minute \(\bf \color{blue}{\cfrac{\frac{50}{3}mile}{5min}\implies \cfrac{50}{3}mile\cdot \cfrac{1}{5min}\implies \cfrac{50mile}{15min}}\qquad thus\\ \quad \\ \cfrac{50\cancel{mile}}{15\cancel{min}}\cdot \cfrac{5280ft}{\cancel{mile}}\cdot \cfrac{\cancel{min}}{60secs}\)